# Vecchia Gaussian Processes: Probabilistic Properties, Minimax Rates and Methodological Developments

### v3

@@ C.1 Proof of Theorem 2
We first prove the theorem with $s=1$ and abbreviate $\hat{Z}^{\tau,1}$ as $\hat{Z}^{\tau}$. Extending the result for general $s>0$ is straightforward and can be achieved by dividing $\epsilon$ by $s$.
Recall that $\eta(i)=j$ denotes that $X_{i}\in\mathcal{N}_{j}$. For all $X_{i}\in\mathcal{X}_{n}$, define $u_{i}$ as
@@
Since $u_{i}$ is a linear transformation of a centered Gaussian random vector, it has also zero mean. Recall also that in view of Theorem 1, $\mathrm{Var}(u_{i})=\mathbb{E}[Z^{\tau}_{X_{i}}-\mathbb{E}(Z^{\tau}_{X_{i}}|Z^{\tau}_{\mathrm{pa}(X_{i})})]^{2}\asymp\tau^{2\alpha}\gamma^{-2\alpha\eta(i)}$. Furthermore, let $j_{0}$ be the smallest integer satisfying $\gamma^{-\alpha j_{0}}\leq\epsilon$ and define the sequence
@@
The rest of the proof consists of three parts. In the first part, we provide an upper bound the $l_{\infty}$ norm of $\hat{Z}_{X_{i}}^{\tau}$ under the event $\cap_{i=1}^{n}A_{i}$. In the second part, we derive a lower bound for the probability of this event $\cap_{i=1}^{n}A_{i}$. In the final part, we combine the previous results and extend them to general scaling parameter $s>0$.
@@
With a slight abuse of notations, for all $1\leq j\leq\eta(n)$, let $\hat{Z}^{\tau}_{\cup_{j^{\prime}=0}^{j}\mathcal{N}_{j^{\prime}}}$ denote the restriction of $\hat{Z}^{\tau}$ to the co-domain $\cup_{j^{\prime}=0}^{j}\mathcal{N}_{j^{\prime}}$. Then recalling the definition of the operator $G_{j}$ in (33), we have for all $0\leq j\leq\eta(i)-1$,
@@
Therefore, by induction of the previous formula, and in view of triangle inequality, the definition of the event $A_{i}$, and Lemma 7, we get
@@
$\lesssim$ $\tilde{\vartheta}_{n}\sum_{j=0}^{\eta(n)}b_{j}\leq 2\tilde{\vartheta}_{n}\tau^{\alpha}\frac{\gamma^{-\alpha j_{0}}}{1-\gamma^{-\alpha/2}}\lesssim\tilde{\vartheta}_{n}\tau^{\alpha}\epsilon.$ (93)
Part 2: We first prove by induction that $u_{1},u_{2},\cdots,u_{n}$ are independent random variables. A single random variable $u_{1}$ is trivially independent. Then, assuming that $u_{1},u_{2},\cdots,u_{l}$ are independent we show below that $u_{l+1}$ is independent from $(u_{1},u_{2},\cdots,u_{l})$. By definition of $u_{i},1\leq i\leq l+1$, we have
@@
which implies independence in view that $u_{i}$s are centered Gaussian random variables. Therefore,
$\mathrm{Pr}(\cap_{i=1}^{n}A_{i})=\prod_{i=1}^{n}\mathrm{Pr}\Big{(}|Z|\leq b_{\eta(i)}/\sqrt{\mathrm{Var}(u_{i})}\Big{)}=\prod_{j=1}^{\eta(n)}\left[\mathrm{Pr}(|Z|\leq b_{j}\tau^{-\alpha}\gamma^{\alpha j})\right]^{|\mathcal{N}_{j}|},$
where $Z\sim N(0,1)$. Next recall some standard tail bounds for the standard normal distribution, i.e.
$\mathrm{Pr}(|Z|\leq u)$ $\geq u/3,\quad\forall 0\leq u\leq 1,$
$\mathrm{Pr}(|Z|\leq u)$ $\geq 1-e^{-u^{2}}\geq\exp\big{(}-2e^{-u^{2}/2}\big{)},\quad\forall\;u\geq 1,$
see (7.2) and (7.3) in [33], respectively. Then, in view of the bound $b_{j}\tau^{-\alpha}\gamma^{\alpha j}\leq 1$ if and only if $j\leq j_{0}$ and noting that (45) implies $|\mathcal{N}_{j}|\lesssim\gamma^{dj}$, we get that for some constant $c_{d}>0$
$\mathrm{Pr}(\cap_{i=1}^{n}A_{i})\geq$ $\prod_{j=1}^{j_{0}}\big{(}\gamma^{-\frac{3}{2}\alpha(j_{0}-j)}/3\big{)}^{c_{d}\gamma^{dj}}\prod_{j=j_{0}+1}^{\eta(n)}\exp\big{(}-2c_{d}\gamma^{dj}e^{-\gamma^{\alpha(j-j_{0})}/2}\big{)}.$ (94)
The logarithm of the first term in (94) can be bounded as
@@
The logarithm of the second term in (94) can be bounded as
@@
By combining the previous three displays and recalling that $\gamma^{-\alpha j_{0}}\asymp\epsilon$, we have get
$\ln\mathrm{Pr}(\cap_{i=1}^{n}A_{i})\lesssim\epsilon^{-d/\alpha}.\\$ (95)
Conclusion: In view of assertions (C.1) and (95),
@@ C.2 Proof of Lemma 8
Let $f\in\mathbb{H}$ be an arbitrary function satisfying $\bm{f}(A)=\bm{f}_{0}(A)$. Then $f(x)-\sum_{i=1}^{|A|}a_{i}K(x_{i},x)=0$, $\forall x\in A$. Then, in view of the reproducing property
@@
Therefore, $K(x,\cdot)$ is orthogonal to $f(\cdot)-\sum_{i=1}^{|A|}a_{i}K(x_{i},\cdot)$ in $\mathbb{H}$, $\forall x\in A$. Thus
@@ C.3 Proof of Lemma 10
In view of equation (2.2), without loss of generality, we can assume that $s=1$ and simplify the notation $Z^{\tau,1}$ as $Z^{\tau}$ and $\mathbb{H}^{\tau,1}$ as $\mathbb{H}^{\tau}$. The proof consists of three steps. In the first step, we control the interpolation error for constant functions. Then in the second step, we prove the lemma for $\beta\in(0,1]$. In the final step, we extend the proof for $\beta>1$.
Let us introduce the notation $B_{i}=\{X_{i}\}\cup\mathrm{pa}(X_{i})$. Then, in view of Condition 2, there exists a $d$-dimensional cube $\mathcal{C}\subset\mathcal{X}$ of length size $c_{L}\gamma^{-\eta(i)}$, which contains $B_{i}$. Also recall Condition 1, that the minimal $\ell^{\infty}$ distance among elements in $\cup_{j^{\prime}=0}^{j}\mathcal{N}_{j^{\prime}}$ is $c_{d}\gamma^{-j}$. Since the cube $\mathcal{C}$ contains $X_{i}$ and its parent set, we have $c_{d}\leq c_{L}$.
@@
$\left|c-\mathbb{E}[Z^{\tau}_{X_{i}}|Z^{\tau}_{\mathrm{pa}(X_{i})}=c\mathbbm{1}]\right|\lesssim\tau^{\alpha}\gamma^{-\eta(i)\alpha},$ (96)
@@
$\inf_{f\in\mathbb{H}^{\tau},\bm{f}(B_{i})=\bm{f}_{0}(B_{i})}\|f\|_{\mathbb{H}^{\tau}}^{2}=\bm{f}_{0}(B_{i})^{T}K_{\tau B_{i},\tau B_{i}}^{-1}\bm{f}_{0}(B_{i})=\inf_{f\in\mathbb{H},\bm{f}(\tau B_{i})=\bm{f}_{0}(B_{i})}\|f\|_{\mathbb{H}}^{2}.$ (97)
Then let us consider a function $\varphi_{0}(\cdot):\mathbb{R}\to\mathbb{R}$ such that $\varphi_{0}\in C^{\alpha+d/2}(\mathbb{R})_{1}$, $\varphi_{0}$ is supported on [2,2][-2,2] and equals a constant on [1,1][-1,1] and define the function $\varphi_{d}(\cdot):\mathbb{R}^{d}\to\mathbb{R}$ as $\varphi_{d}(x)=\prod_{l=1}^{d}\varphi_{0}(x[l])$. Noting that $\varphi_{d}\in C^{\alpha+d/2}(\mathbb{R}^{d})_{1}$ is supported on $2,2]^{d}$ and equals a constant on $1,1]^{d}$, $\varphi_{d}$ also has finite Sobolev norm, i.e., $\|f\|_{W^{\alpha+d/2}}\lesssim 1$. We then divide the proof into two cases based on whether $\tau\gamma^{-\eta(i)}\geq 1$.
If $\tau\gamma^{-\eta(i)}\leq 1$, consider the function
@@
Note that $\tilde{f}(x)=c$ for all $x\in\tau\mathcal{C}$. In view of the proof of Lemma 11.36 of [22], the RKHS norm of the Matérn process is equivalent to the Sobolev norm. Therefore, by recalling (97),
$\inf_{f\in\mathbb{H}^{\tau},\,f(x)=c,\,x\in B_{i}}\|f\|_{\mathbb{H}^{\tau}}^{2}\leq$ $\|\tilde{f}\|_{\mathbb{H}}^{2}\asymp\|\tilde{f}\|_{W^{\alpha+d/2}}^{2}\asymp\|\varphi_{d}\|_{W^{\alpha+d/2}}^{2}\lesssim 1.$ (98)
If $\tau\gamma^{-\eta(i)}\geq 1$, then $\tau\geq 1$. Consider the function
@@
$\inf_{f\in\mathbb{H}^{\tau},\,f(x)=c,\,x\in B_{i}}\|f\|_{\mathbb{H}^{\tau}}^{2}\leq$ $\|\tilde{f}\|_{\mathbb{H}}^{2}\asymp\|\tilde{f}\|_{W^{\alpha+d/2}}^{2}\lesssim(m+1)\|\varphi_{d}\|_{W^{\alpha+d/2}}^{2}\lesssim 1.$ (99)
Plugging in the bounds (98) and (99) into Lemma 9 and using Theorem 1, we get
@@
$\lesssim\sqrt{\tau^{2\alpha}\gamma^{-2\eta(i)\alpha}\cdot 1}=\tau^{\alpha}\gamma^{-\eta(i)\alpha}.$ (100)
@@
$|f_{0}(w)-c|\leq\|w-X_{i}\|_{2}^{\beta}\lesssim\gamma^{-\eta(i)\beta},\;\;\forall w\in B_{i}.$ (101)
Let us introduce the notation $f_{c}(x)=f_{0}(x)-c$, then $|f_{c}(x)|\lesssim\gamma^{-\eta(i)\beta}$, for all $x\in B_{i}$. Note, that assertion (97) holds with $f_{c}$ in place of $f_{0}$. We proceed next by splitting the proof based on the value of $\tau.$
@@
Noting that the minimal separation distance among elements in $B_{i}$ is $c_{d}\gamma^{-\eta(i)}$ and $\forall w\in B_{i},\;\tilde{f}_{w}$ is supported on a cube with side length $c_{d}\tau\gamma^{-\eta(i)}/2$, the supports of $\tilde{f}_{w},w\in B_{i}$ are disjoint. Since $\tilde{f}_{w}(\tau w)=f_{c}(w)$, denoting $\hat{\varphi}_{d}(\cdot)$ as the Fourier transform of $\varphi_{d}(\cdot)$, we have by equation (96),
@@
$\qquad=\tau^{-2\alpha}\gamma^{2\eta(i)(\alpha-\beta)}.$ (102)
@@
By the same analysis as in Step 1, we have
@@
$\asymp\sum_{w\in B_{i}}[f_{c}(w)]^{2}\lesssim\gamma^{-2\eta(i)\beta},$ (103)
where the last step follows from (101). Plugging in the bounds (C.3) and (C.3) into Lemma 9,
@@
$\lesssim$ $\gamma^{-\eta(i)\beta}+\tau^{\alpha}\gamma^{-\eta(i)(\alpha+\beta)}.$ (104)
Then in view of (C.3) and (96), the linearity of the expectation on the condition, and $\gamma^{\eta(i)\beta}\geq 1$,
@@
$\lesssim\tau^{\alpha}\gamma^{-\eta(i)\alpha}+\gamma^{-\eta(i)\beta},$ (105)
@@
Define the operator $\mathcal{L}$ such that $\forall f:\mathcal{X}\to\mathbb{R},\;x\in\mathcal{X}$ , $\mathcal{L}(f)(x)=f(x)-\mathbb{E}[Z^{\tau}_{x}|Z^{\tau}_{\mathrm{pa}(X_{i})}=\bm{f}(\mathrm{pa}(X_{i}))]$. Note that $\mathcal{L}$ is a linear operator and can exchange orders with summation and integration (as long as integration is finite). Then, in view of the induction assumption and inequality (105), for all $x\in\mathcal{C}$

### v4

@@ F.1 Proof of Theorem 2
First, we prove the theorem for $s=1$ and abbreviate $\hat{Z}^{\tau,1}$ as $\hat{Z}^{\tau}$. Extending the result to a general $s>0$ is straightforward and can be accomplished by dividing $\epsilon$ by $s$.
Recall that $\eta(i)=j$ indicates that $X_{i}\in\mathcal{N}_{j}$. For every $X_{i}\in\mathcal{X}_{n}$, define $u_{i}$ as
@@
Since $u_{i}$ is a linear transformation of a centered Gaussian random vector, it also has zero mean. Recall that in view of Theorem 1, $\mathrm{Var}(u_{i})=\mathbb{E}[Z^{\tau}_{X_{i}}-\mathbb{E}(Z^{\tau}_{X_{i}}|Z^{\tau}_{\mathrm{pa}(X_{i})})]^{2}\asymp\tau^{2\alpha}\gamma^{-2\alpha\eta(i)}$. Furthermore, let $j_{0}$ be the smallest integer satisfying $\gamma^{-\alpha j_{0}}\leq\epsilon$ and define the sequence
@@
The remainder of the proof consists of three parts. In the first part, we provide an upper bound for the $L_{\infty}$ norm of $\hat{Z}_{X_{i}}^{\tau}$ under the event $\cap_{i=1}^{n}A_{i}$. In the second part, we derive a lower bound for the probability of this event. In the final part, we combine the previous results and extend them to general scaling parameter $s>0$.
@@
With a slight abuse of notation, for all $1\leq j\leq\eta(n)$, let $\hat{Z}^{\tau}_{\cup_{j^{\prime}=0}^{j}\mathcal{N}_{j^{\prime}}}$ denote the restriction of $\hat{Z}^{\tau}$ to the co-domain $\cup_{j^{\prime}=0}^{j}\mathcal{N}_{j^{\prime}}$. Then, in view of (28), we have for all $0\leq j\leq\eta(i)-1$,
@@
Therefore, by induction of the previous formula, and applying the triangle inequality, the definition of the event $A_{i}$, and Lemma 4, we get
@@
$\lesssim$ $\tilde{\vartheta}_{n}\sum_{j=0}^{\eta(n)}b_{j}\leq 2\tilde{\vartheta}_{n}\tau^{\alpha}\frac{\gamma^{-\alpha j_{0}}}{1-\gamma^{-\alpha/2}}\lesssim\tilde{\vartheta}_{n}\tau^{\alpha}\epsilon.$ (F.1)
Part 2: We first prove by induction that $u_{1},u_{2},\cdots,u_{n}$ are independent random variables. A single random variable $u_{1}$ is trivially independent. Next, assuming that $u_{1},u_{2},\cdots,u_{l}$ are independent, we show that $u_{l+1}$ is independent of $(u_{1},u_{2},\cdots,u_{l})$. By the definition of $u_{i}$, for $1\leq i\leq l+1$, we have
@@
which implies independence, since the $u_{i}$s are centered Gaussian random variables. Therefore,
$\mathrm{Pr}(\cap_{i=1}^{n}A_{i})=\prod_{i=1}^{n}\mathrm{Pr}\Big(|U|\leq b_{\eta(i)}/\sqrt{\mathrm{Var}(u_{i})}\Big)=\prod_{j=1}^{\eta(n)}\left[\mathrm{Pr}(|U|\leq b_{j}\tau^{-\alpha}\gamma^{\alpha j})\right]^{|\mathcal{N}_{j}|},$
where $U\sim N(0,1)$.
Next, recall some standard tail bounds for the standard normal distribution, i.e.
$\mathrm{Pr}(|U|\leq u)$ $\geq u/3,\quad\forall 0\leq u\leq 1,$
$\mathrm{Pr}(|U|\leq u)$ $\geq 1-e^{-u^{2}}\geq\exp\big(-2e^{-u^{2}/2}\big),\quad\forall\;u\geq 1,$
see (7.2) and (7.3) in [34], respectively. Then, noting that $b_{j}\tau^{-\alpha}\gamma^{\alpha j}\leq 1$ if and only if $j\leq j_{0}$ and that (C.1) implies $|\mathcal{N}_{j}|\lesssim\gamma^{dj}$, we obtain, for some constant $c_{d}>0$
$\mathrm{Pr}(\cap_{i=1}^{n}A_{i})\geq$ $\prod_{j=1}^{j_{0}}\big(\gamma^{-\frac{3}{2}\alpha(j_{0}-j)}/3\big)^{c_{d}\gamma^{dj}}\prod_{j=j_{0}+1}^{\eta(n)}\exp\big(-2c_{d}\gamma^{dj}e^{-\gamma^{\alpha(j-j_{0})}/2}\big).$ (F.2)
The logarithm of the first term in (F.2) can be bounded as
@@
The logarithm of the second term in (F.2) can be bounded as
@@
By combining the previous three displays and recalling that $\gamma^{-\alpha j_{0}}\asymp\epsilon$, we obtain
$\ln\mathrm{Pr}(\cap_{i=1}^{n}A_{i})\lesssim\epsilon^{-d/\alpha}.\\$ (F.3)
Conclusion: In view of assertions (F.1) and (F.3),
@@ F.2 Proof of Lemma 5
Let $f\in\mathbb{H}$ be an arbitrary function satisfying $\bm{f}(A)=\bm{f}_{0}(A)$. Then, $f(x)-\sum_{i=1}^{|A|}a_{i}K(x_{i},x)=0$, $\forall x\in A$. In turn, the reproducing property implies that
@@
Therefore, $K(x,\cdot)$ is orthogonal to $f(\cdot)-\sum_{i=1}^{|A|}a_{i}K(x_{i},\cdot)$ in $\mathbb{H}$, $\forall x\in A$. Thus,
@@ F.3 Proof of Lemma 7
In view of equation (2.2), without loss of generality, we can assume $s=1$ and simplify the notation by writing $Z^{\tau,1}$ as $Z^{\tau}$ and $\mathbb{H}^{\tau,1}$ as $\mathbb{H}^{\tau}$. The proof consists of three steps. In the first step, we control the interpolation error for constant functions. In the second step, we prove the lemma for $\beta\in(0,1]$. In the final step, we extend the proof to $\beta>1$.
Let us introduce the notation $B_{i}=\{X_{i}\}\cup\mathrm{pa}(X_{i})$. By Condition 2, there exists a $d$-dimensional cube $\mathcal{C}\subset\mathcal{X}$ of side length $c_{L}\gamma^{-\eta(i)}$ that contains $B_{i}$. Also, by Condition 1, the minimal $L_{\infty}$ distance among elements in $\cup_{j^{\prime}=0}^{j}\mathcal{N}_{j^{\prime}}$ is $c_{d}\gamma^{-j}$. Since the cube $\mathcal{C}$ contains $X_{i}$ and its parent set, we have $c_{d}\leq c_{L}$.
@@
$\left|c-\mathbb{E}[Z^{\tau}_{X_{i}}|Z^{\tau}_{\mathrm{pa}(X_{i})}=c\mathbbm{1}]\right|\lesssim\tau^{\alpha}\gamma^{-\eta(i)\alpha},$ (F.4)
@@
$\inf_{f\in\mathbb{H}^{\tau},\bm{f}(B_{i})=\bm{f}_{0}(B_{i})}\|f\|_{\mathbb{H}^{\tau}}^{2}=\bm{f}_{0}(B_{i})^{T}K_{\tau B_{i},\tau B_{i}}^{-1}\bm{f}_{0}(B_{i})=\inf_{f\in\mathbb{H},\bm{f}(\tau B_{i})=\bm{f}_{0}(B_{i})}\|f\|_{\mathbb{H}}^{2}.$ (F.5)
Next, let us consider a function $\varphi_{0}(\cdot):\mathbb{R}\to\mathbb{R}$ such that $\varphi_{0}\in C^{\alpha+d/2}(\mathbb{R})_{1}$, $\varphi_{0}$ is supported on [2,2][-2,2] and equals to a constant on [1,1][-1,1]. We define the function $\varphi_{d}(\cdot):\mathbb{R}^{d}\to\mathbb{R}$ as $\varphi_{d}(x)=\prod_{l=1}^{d}\varphi_{0}(x[l])$. Note that $\varphi_{d}\in C^{\alpha+d/2}(\mathbb{R}^{d})_{1}$ is supported on $2,2]^{d}$, equals to a constant on $1,1]^{d}$, and has finite Sobolev norm $\|f\|_{W^{\alpha+d/2}}\lesssim 1$. We then split the proof into two cases based on whether $\tau\gamma^{-\eta(i)}\geq 1$.
If $\tau\gamma^{-\eta(i)}\leq 1$, take the function
@@
Note that $\tilde{f}(x)=c$ for all $x\in\tau\mathcal{C}$. In view of the proof of Lemma 11.36 of [22], the RKHS norm of the Matérn process is equivalent to the Sobolev norm. Therefore, by recalling (F.5),
$\inf_{f\in\mathbb{H}^{\tau},\,f(x)=c,\,x\in B_{i}}\|f\|_{\mathbb{H}^{\tau}}^{2}\leq$ $\|\tilde{f}\|_{\mathbb{H}}^{2}\asymp\|\tilde{f}\|_{W^{\alpha+d/2}}^{2}\asymp\|\varphi_{d}\|_{W^{\alpha+d/2}}^{2}\lesssim 1.$ (F.6)
If $\tau\gamma^{-\eta(i)}\geq 1$, then $\tau\geq 1$. Take the function
@@
$\inf_{f\in\mathbb{H}^{\tau},\,f(x)=c,\,x\in B_{i}}\|f\|_{\mathbb{H}^{\tau}}^{2}\leq$ $\|\tilde{f}\|_{\mathbb{H}}^{2}\asymp\|\tilde{f}\|_{W^{\alpha+d/2}}^{2}\lesssim(m+1)\|\varphi_{d}\|_{W^{\alpha+d/2}}^{2}\lesssim 1.$ (F.7)
Plugging in the bounds (F.6) and (F.7) into Lemma 6 and using Theorem 1, we get
@@
$\lesssim\sqrt{\tau^{2\alpha}\gamma^{-2\eta(i)\alpha}\cdot 1}=\tau^{\alpha}\gamma^{-\eta(i)\alpha}.$ (F.8)
@@
$|f_{0}(w)-c|\leq\|w-X_{i}\|_{2}^{\beta}\lesssim\gamma^{-\eta(i)\beta},\;\;\forall w\in B_{i}.$ (F.9)
Let us introduce the notation $f_{c}(x)=f_{0}(x)-c$. Then, $|f_{c}(x)|\lesssim\gamma^{-\eta(i)\beta}$, for all $x\in B_{i}$. Note that assertion (F.5) holds with $f_{c}$ in place of $f_{0}$. We next proceed by splitting the proof based on the value of $\tau.$
@@
Noting that the minimal separation distance among elements in $B_{i}$ is $c_{d}\gamma^{-\eta(i)}$ and $\forall w\in B_{i},\;\tilde{f}_{w}$ is supported on a cube with side length $c_{d}\tau\gamma^{-\eta(i)}/2$, the supports of $\tilde{f}_{w},w\in B_{i}$ are disjoint. Since $\tilde{f}_{w}(\tau w)=f_{c}(w)$, denoting $\hat{\varphi}_{d}(\cdot)$ as the Fourier transform of $\varphi_{d}(\cdot)$, we have by equation (F.4),
@@
$\qquad=\tau^{-2\alpha}\gamma^{2\eta(i)(\alpha-\beta)}.$ (F.10)
@@
By the same arguments as in Step 1,
@@
$\asymp\sum_{w\in B_{i}}[f_{c}(w)]^{2}\lesssim\gamma^{-2\eta(i)\beta},$ (F.11)
where the last step follows from (F.9). Plugging in the bounds (F.3) and (F.3) into Lemma 6, yields
@@
$\lesssim$ $\gamma^{-\eta(i)\beta}+\tau^{\alpha}\gamma^{-\eta(i)(\alpha+\beta)}.$ (F.12)
Then in view of (F.3) and (F.4), the linearity of the expectation on the condition, and $\gamma^{\eta(i)\beta}\geq 1$,
@@
$\lesssim\tau^{\alpha}\gamma^{-\eta(i)\alpha}+\gamma^{-\eta(i)\beta},$ (F.13)
@@
Define the operator $\mathcal{L}$ such that $\forall f:\mathcal{X}\to\mathbb{R},\;x\in\mathcal{X}$ , $\mathcal{L}(f)(x)=f(x)-\mathbb{E}[Z^{\tau}_{x}|Z^{\tau}_{\mathrm{pa}(X_{i})}=\bm{f}(\mathrm{pa}(X_{i}))]$. Note that $\mathcal{L}$ is a linear operator and can exchange orders with summation and integration (as long as integration is finite). Then, in view of the induction assumption and inequality (F.13), for all $x\in\mathcal{C}$,